Circular No. 8135

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 K1 (SPACEWATCH) An object initially reported as asteroidal by J. A. Larsen on CCD

images obtained with the 0.9-m Spacewatch reflector (discovery observation below) was posted on the NEO Confirmation Page. CCD images taken by A. Lopez and R. Pacheco (Mallorca, 0.41-m reflector) on May 23.9 UT showed cometary appearance (and m_{1} = 18.2–18.6). A. E. Gleason found the coma to be quite obvious on May 24.3 images taken with the 1.8-m Spacewatch II reflector at Kitt Peak, and Larsen found a 10′′ coma on Spacewatch I images taken on May 24.4.

2003 UT  α_{2000}  δ_{2000}  m_{2}
May 23.37801  16^{h}11^{m}27.^{s}91  −5^{o}28′12′′.7  20.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-K37.

T = 2003 Jan. 10.761 TT  ω = 327.^{o}390
Ω = 250.809 }2000.0
q = 2.34873 AU  i = 131.957

COMET P/2003 H4 (LINEAR) The following orbital elements, from MPEC 2003-K34, indicate that

this comet passed 0.07 AU from Jupiter in June 1929, before which q and P were larger.

T = 2003 May 14.3119 TT  ω = 10.^{o}4498
e = 0.490221  Ω = 226.7962 }2000.0
q = 1.703366 AU  i = 18.1550
a = 3.341379 AU n^{o} = 0.1613673 P = 6.11 years

SUPERNOVA 2003el IN NGC 5000 Further to IAUC 8134, M. Papenkova and W. Li report the LOTOSS

discovery of an apparent supernova on unfiltered KAIT images taken on May 22.3 (mag ∼ 18.8) and 23.3 UT (mag ∼ 18.9). SN 2003el is located at α = 13^{h}09^{m}46^{s}.20, δ = +28^{o}54′21′′.6 (equinox 2000.0), which is 16′′.7 west and 2′′.7 south of the nucleus of NGC 5000. KAIT images taken on Apr. 19.3 (limiting mag ∼ 19.0), and Mar. 7.3 (limiting mag ∼ 19.6) showed nothing at this position.

2003 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green